# ICONS: Influence Consensus for Vision-Language Data Selection

### v2

@@ 3.2 Comparisons with Baselines
Figure 3: Data Overlap between Specialists and Generalist Selection. It shows varying degrees of overlap, ranging from 3.27% (VQAv2 [9]) to 24.21% (LLAVA-W Bench [23]).
Figure 4: Performance Comparison at Different Selection Ratios. Here we show the average relative performance comparisons of various data selection methods at different selection ratios. Our ICONS consistently outperforms all baseline approaches across different selection ratios, achieving 99.2% relative performance with 40% of the data, and remarkably exceeding 102% at 60% selection ratio. This shows that our selected subset is not only more efficient but can be even more effective than the full dataset for model training.
3.3 Analysis
From Specialist to Generalist.  In order to understand the effectiveness of the intermediate task-specific influence matrices we obtained from the specialist stage, we further conduct experiments on selecting 20% of data for individual tasks. As shown in Tab. 3, using 20% task-specific data, it achieves comparable or even superior performance to full data training in several cases (e.g., MME from 1476.9 to 1506.1 and SQA-I from 68.4 to 69.8). Then, through our influence consensus mechanism in the generalist stage, we identify consistently influential cross-tasks samples and select a 20% subset. This generalist approach maintains remarkably strong performance, with an average drop of 1.33% across most tasks compared to specialist baselines.
As visualized in Fig. 3, the overlap between specialist and generalist selections varies significantly across tasks, ranging from minimal overlap in simpler tasks like VQAv2 [9] (3.27%) and VizWiz [10] (3.28%) to substantial agreement in more complex tasks like LLAVA-W Bench [23] (24.21%). Notably, some tasks even show performance improvements under the generalist selection, such as SQA-I [25] (improving by 1.43%) and POPE [21] (improving by 1.04%). This empirically validates our consensus-based approach: by first understanding task-specific influence patterns and then building consensus across tasks, we can effectively identify a compact, universal training set that maintains strong performance across diverse vision-language tasks while significantly reducing the required training data.
Different Selection Ratio.  We evaluate our ICONS against baseline methods across different selection ratios, ranging from 5% to 60% of the training data. As shown in Fig. 4, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial. Another notable observation is that as the selection ratio increases, the performance gap between different methods gradually narrows. This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information. Despite this convergence trend, ICONS consistently maintains its performance edge across all selection ratios, and remarkably reaching above 100% at 60% selection ratio. One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
(a) Pairwise Overlap Heatmap between Specialist Benchmarks. The values are the percentage of selected overlapping samples across different benchmarks, with high overlap seen between POPE and GQA (60.2%) and between VQAv2 and VizWiz (49.0%). Lower overlap, such as 3.3% between MMBench(en) and GQA, shows more distinct selected sample sets.
(b) Pairwise Overlap Heatmap between Different Aggregation Approaches. The values are the percentage of overlapping selected samples between each pair of aggregation approaches. Our Vote-based influence consensus approach shows moderate overlap (53.8%-70.5%) with other methods, while Mean and Max selections are highly correlated (99.9%).
Figure 5: Pairwise Overlap Heatmaps between Different Aggregation Approaches and Benchmarks. (a) compares the overlap between different specialist benchmarks, while (b) shows overlap between various aggregation methods.
@@ 3.3 Analysis
Divergent Cross-Task Influence Patterns.  As shown in Fig. 5(a), the pairwise overlap heatmap reveals significant variation in training data influence rankings across different tasks. High overlap, such as between VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks. However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others. Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples. These findings empirically demonstrate significant overlap in influential samples across tasks, motivating the use of influence consensus for effective multi-task data selection.
3.4 Transferability
@@
We evaluate the transferability of our data selection approach across two dimensions: unseen-task generalization and cross-architecture-scale generalization. Unseen-task Generalization.  LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not considered during our data selection process. As shown in Tab. 4, we test across a diverse spectrum of tasks including MMVet [45], NaturalBench [19], AI2D [16], ChartQA [29], DocVQA [30], InfoVQA [31], RealWorldQA [41] and CMMMU [46]. LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training. Notably, we observe improvements on InfoVQA (103.8%), NaturalBench (105.5%), RealWorldQA (104.4%), and CMMMU (113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process. Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks - for example, improving performance on ChartQA (17.1 vs. 15.1), MMVet (29.7 vs. 27.6), CMMMU (25.2 vs. 21.9) and NaturalBench (12.8 vs. 11.1). This consistent improvement over random selection further validates the effectiveness of ICONS. This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains, and consistently outperforms random selection by a significant margin.
@@
Table 4: Unseen-Task Generalization. Performance comparison on unseen benchmarks when trained on the selected subset. The relative performance ratios show that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains. Notably, we observe improvements on InfoVQA [31] (103.8%), MMVet [45] (95.5%), NaturalBench [19] (103.2%), RealWorldQA [41] (104.4%), and CMMMU [46] (114.0%), suggesting that in some cases, training on our LLaVA-ICONS-133K may even outperform using the full dataset.
Cross-Architecture-Scale Generalization.  We further conduct experiments on cross architecture scale generalization to evaluate the transferability of our selected data across different model scales. While our subset was initially selected using LLaVA-1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-1.5-13B. This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns. Our results in Tab. 5 show cross architecture scale generalization, with the 7B-selected achieving 98.1% relative performance, suggesting that our selected subset captures fundamental visual-language understanding capabilities that scale effectively across model sizes.
@@
VQAv2 GQA Vizwiz SQA-I TextVQA POPE MME MMBench MMBench(cn) LLAVA-W Rel. (%)
@@
Table 5: Cross-Architecture-Scale Generalization. Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7b-lora (7B-selected) shows strong cross-architecture transferability, achieving competitive performance when trained with data selected via LLaVA-v1.5-13b-lora (13B-selected). Specifically, while the 7B-selected achieves 97.3% relative performance, the 13B-selected model reaches 98.1%, showing that our selected subset’s effectiveness scales well to larger architectures. Both significantly outperform random selection (95.7%), with the 13B-selected option showing particular strength in complex reasoning tasks like MMBench and POPE. This consistency across model scales highlights that our data selection approach captures fundamental visual-language understanding patterns that generalize well across different model architecture scales.
3.5 Ablation Studies
Different Aggregation Approaches.  To understand how different methods of combining task-specific influence scores affect the performance, we compare our majority voting approach (Vote) with four aggregation approaches: (1) Mean, which averages scores across tasks, (2) Max, which selects the highest score per sample, (3) Rank, which averages the relative ranking of samples within each task, (4) Norm, which normalize the scores before averaging:
$I_{\text{mean}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}I_{k}({\bm{z}}_{i})\qquad I_{\text{norm}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
$I_{\text{max}}({\bm{z}}_{i})=\max_{k=1,\ldots,K}I_{k}({\bm{z}}_{i})\qquad I_{\text{rank}}({\bm{z}}_{i})=\frac{1}{K}\sum_{k=1}^{K}\text{rank}(I_{k}({\bm{z}}_{i}))$
Results in Tab. 6 show that our Vote consistently outperforms alternatives across most tasks, achieving the highest overall relative performance (Rel.) of 98.6%. While methods like Rank perform well for MME (1490.0 vs. 1485.7 with Norm), they underperform overall. To better understand the relationships between these aggregation approaches, we visualize their pairwise selection overlap in Fig. 5(b). Specifically, the Mean and Max approaches show an extremely high overlap of 99.9%, suggesting they tend to select very similar influential samples. In contrast, our Vote-based approach exhibits a more moderate overlap with other methods, ranging from 53.8% to 70.5%, potentially those that are influential across a broader variety of tasks rather than overfitting to specific tasks.
Task Mean Max Rank Norm Vote (ours)
VQAv2 [9] 75.7 75.2 74.9 75.1 76.3
GQA [13] 59.6 59.8 58.6 60.1 60.7
VizWiz [10] 47.9 48.1 40.5 46.4 50.1
SQA-I [25] 65.5 66.2 69.8 69.8 70.8
TextVQA [37] 55.5 55.5 55.2 54.5 55.6
POPE [21] 86.0 85.5 85.6 85.6 87.5
MME [7] 1422.1 1470.7 1490.0 1482.6 1485.7
MMBench (en) [47] 59.0 58.3 59.0 58.9 63.1
MMBench (cn) [47] 51.0 51.8 50.8 52.5 55.8
LLaVA-W Bench [23] 66.2 66.2 66.4 66.3 66.1
Rel.(%) 96.4 96.1 95.9 96.8 98.6
Table 6: Comparison of Different Aggregation Approaches. Performance comparison of different methods for combining task-specific influence scores, using 20% of training data. All scores are relative to full dataset training.
Task Full Direct Consensus Delta (%)
merge aware (ours)
VQAv2 [9] 79.1 76.1 76.3 0.26
GQA [13] 63.0 59.4 60.7 2.19
VizWiz [10] 47.8 46.1 50.1 8.67
SQA-I [25] 68.4 68.7 70.8 3.06
TextVQA [37] 58.2 54.1 55.6 2.77
POPE [21] 86.4 85.1 87.5 2.82
MME [7] 1476.9 1419.2 1485.7 4.69
MMBench (en) [47] 66.1 61.9 63.1 1.94
MMBench (cn) [47] 58.9 50.3 55.8 10.94
LLaVA-W Bench [23] 67.9 65.2 66.1 1.38
Rel. (%) 100 94.7 98.6 -
Table 7: Comparison of Consensus-aware vs. Direct-merge Selection. Performance comparison across different benchmarks using 20% of training data. Direct-merge computes influence directly on combined validation sets, while consensus-aware refers to our specialist-to-generalist approach. The Delta (%) column represents the relative improvement from Direct-merge to Consensus-aware. Consensus-aware outperforms Direct-merge across all benchmarks, while Direct-merge notably underperforms even the random baseline.
Consensus-aware Selection vs. Direct-merge Selection.  To validate our choice of a consensus-aware approach (first computing task-specific influence matrices and then aggregating results) versus a direct-merge approach (directly computing influence using a combined validation set), we compare the performance of both approaches. In the direct-merge approach, we combine validation sets from all tasks into a single pool, denoted as $\{{\bm{z}}^{\prime}\}_{\mathrm{merged}}$, and compute influence scores directly against this combined set: $\mathcal{I}_{\text{direct}}({\bm{z}}\rightarrow\{{\bm{z}}^{\prime}\}_{\mathrm{merged}})$. As described in Tab. 7, our consensus-aware approach consistently outperforms direct-merge selection across all benchmarks, with the largest gains observed on complex tasks like MME [7].
A potential reason could be the direct-merge approach does not account for the different vision-language task difficulties or validation set sizes across the target tasks. By pooling all validation examples, it effectively treats each validation data from all tasks equally, which can lead to the influence scores being dominated by tasks with more samples or with higher influence scores on average. The task-aware voting-based influence consensus mechanism of our direct-merge approach are more robust to these issues. Additionally, our approach is more scalable, as it avoids computing a single, large influence matrix across validation data from all target tasks. Adding new tasks only requires computing influence scores for those specific tasks and updating the voting mechanism, while direct-merge selection would need to recompute the entire large influence matrix which makes it less practical.
4 Related Work
Visual Instruction Tuning.
Multimodal large language models (MLLMs), e.g., Flamingo [2], LLaVA [23], BLIP2 [20], and Cambrian [39], enhance the capabilities of large language models (LLMs) on various multimodal tasks. A key component in advancing MLLMs is visual instruction tuning [23], a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants. This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications [23].
Data Selection.
Data selection methods can be categorized based on the types of information they utilize for selection. Representation-based approaches [1, 18] leverage neural embeddings to capture data representations. Loss trajectory-based methods [32] prioritize data points that contribute most significantly to reducing generalization error over training. Gradient-based techniques [33, 42, 6] select data based on gradient information. Recent work has explored various approaches to select optimal visual instruction tuning datasets. Concurrent work TIVE [24] employs gradient-based selection to identify representative instances. TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty. Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks. COINCIDE [18] clusters data based on representations associated with concept-skill compositions. Our work follows targeted instruction tuning selection approach LESS [42] to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
5 Conclusion
In this work, we introduce ICONS, a simple yet effective influence consensus-based approach for visual instruction tuning data selection. Our two-stage specialist-to-generalist method, first computes task-specific influence matrices and then select a subset by prioritizing universally influential samples through a cross-task voting mechanism. Through extensive experiments, we show that ICONS achieving 98.6% of full dataset performance using only 20% of the data and it consistently outperform baseline methods across different selection ratios. Additionally, we release LLaVA-ICONS-133K, a 20% subset of LLaVA-665K dataset, that not only maintains strong performance across diverse vision-language tasks but also shows transferability to unseen tasks. We hope our work inspires further exploration into data-efficient methods for vision-language models across diverse applications.
Limitations.  Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (§B.2). This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets. To support broader research community, we release our LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
Acknowledgments.  This material is based upon work supported by the National Science Foundation under Grant No. 2107048 and No.2112562. Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This work is also supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001). We thank many people for their helpful discussion and feedback, listed in alphabetical order by last name: Allison Chen, Hamish Ivison, Carlos E. Jimenez, Polina Kirichenko, Jaewoo Lee, Tiffany Ling, Zhiqiu Lin, Shengbang Tong, Ethan Tseng, Justin Wang, Zirui Wang.
Appendix for ICONS: Influence Consensus for Vision-Language Data Selection
In this supplement, we provide additional experiment details and ablation studies (§A) and additional analysis (§B). We detail our algorithmic framework (§C) and discuss future research directions (§D). Finally, we include visualization of the selected data (§E) to better understand data samples with high influence.
Appendix A Additional Experiment Details & Ablations
Projection Dimension.  We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions. The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space. Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count. We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig. 6.
Warm-up Ratio.  To initiate training, we use a warm-up set including 5% of the total training data. We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig. 7. Our experiments reveal that increasing the warm-up data size does not lead to performance improvements. Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%). Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%. This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
Figure 6: Projection Dimension Ablation. We show the performance of ICONS at different projected dimensions (1024, 2560, 5120, 10240), compared to the random baseline. The performance increases with the projected dimension and reaches a plateau around dimension 5120.
Figure 7: Warm-up Ratio Ablation. The blue line represents ICONS performance across different warm-up ratios (5%, 10%, 20%, and 100%), while the red dashed line shows the random baseline performance. Results show that smaller warm-up ratios (5-20%) achieve better performance compared to using the full dataset (100%).
Appendix B Additional Analysis
B.1 Consistency Analysis
To evaluate the consistency of our approach, we conduct three independent runs of the experiment. As shown in Fig. 8, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over the random baseline, which achieves 95.8$\pm$2.7%. The lower standard deviation in our results (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.

### v3

@@ 3.3 Analysis of aggregation strategies
Figure 4: Data overlap between specialists and generalist selection. Overlap varies significantly, from 3.27% (VQAv2) to 24.21% (LLAVA-W Bench), reflecting varying alignment between task-specific and consensus selections.
@@
Divergent multi-task influence patterns.  As shown in Fig. 4, the pairwise overlap heatmap shows notable variation in training data influence across tasks. High overlap – e.g., VQAv2 and VizWiz (49.0%) or POPE and GQA (60.2%), suggests that certain samples are beneficial across similar tasks. However, low overlap, like the 3.3% between MMBench (en) and GQA, highlights that highly influential samples for one task may have limited impact on others. Even closely related tasks, such as MMBench in different languages (English and Chinese), share 67.4% of influential samples. To understand task-specific influence matrices from the specialist stage, we select the top 20% samples per task (Specialists). Overlap with our generalist subset (Fig. 4) varies significantly, from minimal in tasks like VQAv2 (3.27%) and VizWiz (3.28%) to substantial agreement in tasks like LLAVA-W Bench llava (24.21%). These findings empirically demonstrate significant overlap in influential samples across tasks and validate our approach: by analyzing task-specific gradient-based influence patterns and building consensus across tasks, we can identify a compact subset that captures broadly useful samples across tasks, yielding strong performance with significantly less data.
3.4 ICONS generalizes to unseen tasks
@@
LLaVA-ICONS-133K demonstrates exceptional generalization on entirely unseen benchmarks that were not used during data selection. As shown in Tab. 4, we test across a diverse spectrum of tasks including MMVet mmvet , NaturalBench naturalbench , AI2D ai2d , ChartQA chartqa , DocVQA docvqa , InfoVQA infographicvqa , RealWorldQA xai2024grok and CMMMU cmmmu . LLaVA-ICONS-133K achieves 95.5-113.9% (Rel.) compared to full dataset training (InfoVQA: 103.8%, NaturalBench: 105.5%, RealWorldQA: 104.4%, and CMMMU: 113.9%), suggesting that, in some cases, training on LLaVA-ICONS-133K may even outperform training on the full dataset, despite these tasks not being included in the selection process. Importantly, LLaVA-ICONS-133K significantly outperforms random selection across all benchmarks. This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains. We further provide the cross-architecture generalization results in Appendix §D.
@@
Table 4: Unseen-task generalization. Performance comparison on unseen benchmarks when trained on selected subsets. Notably, we observe improvements on InfoVQA (103.8%), RealWorldQA (104.4%), and CMMMU (114.0%), highlighting strong generalization to unseen tasks.
3.5 ICONS outperforms baselines across ratios and exceeds full-data training at 60%
Figure 5: Different selection ratios. ICONS consistently outperforms all baselines across different selection ratios and remarkably exceeding 102% at 60% selection ratio.
To understand how ICONS scales, we evaluate it across different selection ratios, progressively scaling the subset size from 5% to 60% of LLaVA-665K. As shown in Fig. 5, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial. Second, as the selection ratio increases, the performance gap between different methods gradually narrows. This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information. Despite this convergence trend, ICONS consistently outperforms all baselines across all selection ratios. Remarkably, it even surpasses full dataset performance at the 60% ratio, achieving over 102% relative score. One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
4 Conclusion
In this work, we introduce ICONS, a simple yet effective influence consensus-based approach for visual instruction tuning data selection. By leveraging gradient-based influence estimation and aggregating task-specific selections through majority voting, our two-stage specialist-to-generalist approach selects training examples that are broadly beneficial across multiple downstream tasks. ICONS addresses limitations of prior selection methods by avoiding assumptions about score comparability across tasks and reducing the sensitivity to outlier task rankings, which can bias selection in both score-based and rank-based approaches. Beyond data selection, it provides a principled way to reason about data influence in multitask mixtures. Through extensive experiments, we show that ICONS builds compact, high-impact datasets without sacrificing performance or generalization, achieving 98.6% of full dataset performance using only 20% of LLaVA-665K and generalizing well to unseen tasks or architectures. We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, 20% subsets of the LLaVA-665K and Cambrian-7M datasets, maintaining strong performance on diverse tasks and transferring well to unseen ones. We hope our work inspires further exploration into data-efficient methods for vision-language models across diverse applications.
Appendix A Computational complexity
A.1 Complexity analysis
Computing gradient-based influence requires a non-trivial amount of computational resources. In the specialist stage, the complexity scales with both the dataset size $|\mathcal{D}|$ and the gradient dimension $d$. This stage consists of three steps. First, the warm-up training has a complexity of $\mathcal{O}(|\mathcal{D}_{\text{warmup}}|)$. Second, the gradient computation stage has a computational complexity of $\mathcal{O}(|\mathcal{D}|+|\mathcal{D}_{\text{val}}|)$ for forward and backward passes, with storage requirements of $\mathcal{O}(|\mathcal{D}|\cdot d+|\mathcal{D}_{\text{val}}|\cdot d)$ for the gradients. Third (and finally), the influence matrix computation requires $\mathcal{O}(|\mathcal{D}|\cdot|\mathcal{D}_{\text{val}}|\cdot d^{\prime})$ compute cost, where $d^{\prime}$ is the reduced dimension after projection. The generalist stage, focusing on influence consensus across tasks, has lower computational requirements. It begins with threshold computation, requiring $\mathcal{O}(K\cdot|\mathcal{D}|\log|\mathcal{D}|)$ operations for sorting across $K$ tasks. The voting process then takes $\mathcal{O}(K\cdot|\mathcal{D}|)$ compute, followed by a final selection step with complexity $\mathcal{O}(|\mathcal{D}|\log|\mathcal{D}|)$ for sorting the aggregated votes. Storage requirements for this stage are minimal, primarily for the final selected subset.
A.2 Resource requirements
A.3 Discussion on cost-benefit justification
Although gradient-based data selection is computationally intensive, we argue that the initial cost is justified by three key considerations. First, the computational expense is largely a one-time investment: once gradients are computed, they can be stored in our gradient datastore and reused across multiple model iterations, target tasks, and diverse downstream applications. This reusability becomes especially valuable as the number of target datasets grows, because each new target dataset can leverage existing gradient computations, making the selection increasingly efficient at scale.
Second, our empirical results demonstrate substantial performance benefits. Training on a strategically chosen 60% subset of data not only reduces training time but also surpasses the performance obtained by using the full dataset. This improvement underscores how directing more compute resources toward a carefully selected subset can yield higher returns on a fixed set of data.
Lastly, the initial compute-intensive investment in data selection is amortized across future training iterations and future developers. By leveraging the curated, higher-quality dataset, they can substantially reduce training costs.
Appendix B Related work
B.1 Visual instruction tuning
Multimodal large language models (MLLMs), e.g., Flamingo flamingo , LLaVA llava , BLIP2 blip-2 , and Cambrian (tong2024cambrian, ), enhance the capabilities of large language models (LLMs) on various multimodal tasks. A key component in advancing MLLMs is visual instruction tuning (llava, ), a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants. This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications (llava, ).
B.2 Data selection
Data selection methods survey can be categorized based on the types of information they utilize for selection. Representation-based approaches (semdedup, ; coincide, ) leverage neural embeddings to capture data representations. Loss trajectory-based methods (mindermann2022prioritized, ) prioritize data points that contribute most significantly to reducing generalization error over training. Gradient-based techniques (el2n, ; xia2024less, ; deng2024influential, ) select data based on gradient information. Recent work has explored various approaches to select optimal visual instruction tuning datasets. Concurrent work TIVE (liu2024less, ) employs gradient-based selection to identify representative instances. TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty. Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks. COINCIDE (coincide, ) clusters data based on representations associated with concept-skill compositions. Our work follows targeted instruction tuning selection approach LESS xia2024less to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information from various tasks and selecting data for multiple downstream tasks via majority voting.
Appendix C Additional experiment details & ablations
Task MME POPE SQA-I MMBench VQAv2 GQA VizWiz TextVQA LLaVA-W
en cn
$|\mathcal{D}_{\text{val}}|$ 986 500 424 1,164 1,164 1,000 398 800 84 84
$|\mathcal{D}_{\text{test}}|$ 2,374 8,910 4,241 1,784 1,784 36,807 12,578 8,000 5,000 84
Task Type Y/N Y/N MCQ MCQ MCQ VQA VQA VQA VQA VQA
C.1 Additional task details
Here, we provide further details on the target tasks, as summarized in Tab. 5. These tasks cover a wide range of multimodal benchmarks commonly used, including Yes/No questions (Y/N), multiple-choice understanding questions (MCQ) and visual question answering (VQA).
C.2 Projection dimension
We primarily set the projection dimension to 5120, reducing features from 338.7M to 5120 dimensions. The choice of 5120 was empirically validated for its trade-off between effective capturing gradient representation and maintaining a manageable parameter space. Our LLaVA-v1.5-7b-lora architecture includes a total of 7.4B parameters, with 338.7M parameters being trainable after LoRA adaptation, accounting for approximately 4.58% of the total parameter count. We further ablate different projection dimensions (1024, 2560, 5120, and 10240), with results provided in Fig. 6(a).
C.3 Warm-up Ratio
To initiate training, we use 5% of the total training data. We conducted ablation studies to evaluate the impact of varying warm-up ratios (5%, 10%, 20%, and 100%) on selection performance, as shown in Fig. 6(b). Our experiments reveal that increasing the warm-up data size does not lead to performance improvements. Surprisingly, models trained with smaller warm-up ratios (5-20%) consistently outperform those trained with the full dataset (100%). Specifically, the 5% warm-up ratio achieves the best performance at 98.6%, while using the complete dataset results in a performance drop to 97.8%. This finding suggests that a small subset of training data is sufficient and even beneficial for model initialization, and potentially gives better signals in the early training stages.
(a) Projection Dimension Ablation. We show the performance of ICONS at different projected dimensions (1024, 2560, 5120, 10240), compared to the random baseline. The performance increases with the projected dimension and reaches a plateau around dimension 5120.
(b) Warm-up Ratio Ablation. The blue line represents ICONS performance across different warm-up ratios (5%, 10%, 20%, and 100%), while the red dashed line shows the random baseline performance. Results show that smaller warm-up ratios (5-20%) achieve better performance compared to using the full dataset (100%).
Figure 6: Ablation studies on (left) projection dimension and (right) warm-up ratio.
Appendix D Additional analysis
In this section, we provide additional analyses to better understand different aspects of our ICONS. We begin by analyzing the effectiveness of task-specific selections and their aggregation into a generalist subset (§D.1). We further explore whether incorporating visual dependency information into the selection process affects performance across different types of vision-language tasks (§D.2). We also evaluate the transferability of our selected subset across different model scales (§D.3). Additionally, we evaluate the consistency of our method across multiple runs (§D.4), showing its robustness and reliability.
@@ Appendix D Additional analysis
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9
Specialist 77.1 61.1 53.1 69.8 55.7 86.6 1506.1 66.0 56.4 67.1
Generalist 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
Delta (%) 1.04 0.65 5.65 -1.43 0.18 -1.04 1.35 4.34 1.06 1.49
Table 6: Single-task Selection (Specialist) vs. Consensus-aware Multi-task Selection (Generalist). The single-task data selection approach selects 20% of LLaVA-665K per task, while our consensus-aware multi-task data selection approach selects a total of 20% data across all tasks.
D.1 From specialist to generalist
To understand the intermediate task-specific influence matrices we obtained from the specialist stage, we select 20% of data for each individual task. The task-specific data (Specialists) achieves comparable or superior performance than full data training (Tab. 6). With influence consensus at the generalist stage, we select a 20% subset with only a 1.33% average drop across tasks compared to specialist baselines. This validates our approach: by understanding task-specific influence patterns and building consensus across tasks, we can identify a compact, universal training set that maintains strong performance with significantly less data.
D.2 Visual dependency influence ranking
Recent work tong2024cambrian has shown that vision-language tasks vary in their reliance on visual information: tasks like MMBench mmbench depend heavily on visual grounding, while others like SQA-I sqa can be handled primarily through language, showing only a 5% drop in performance when visual input is removed tong2024cambrian . To take visual dependency of training data into account, we further explored gradient-based Visual Dependency Score (VDS). For each data point, we calculate the gradient of the model’s auto-regressive cross-entropy loss with both the original image and a Gaussian noise image $I_{\text{noise}}\sim\mathcal{N}(0,1)$, keeping the text input constant. This quantifies how much the visual component contributes to model performance. We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ w.r.t the model’s gradient alignment and visual dependency. $\mathcal{I}_{\text{VDS}}$ is computed as:
Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different vision-language tasks (Tab. 7). The approach shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%). However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%). These results align with and extend the findings in Cambrian tong2024cambrian , demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly. This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in training.
VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLaVA-W Bench
Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9
w/o VDS 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1
w/ VDS 75.8 60.9 50.3 69.5 54.8 86.8 1489.3 64.3 56.3 67.9
Delta (%) -0.66 +0.33 +0.40 -1.84 -1.44 -0.80 +0.24 +1.90 +0.90 +2.72
Table 7: Impact of Visual Dependency Score (VDS) on Selection Performance. Rows show performance without VDS, with VDS, and the performance change (Delta). VDS improves performance on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), but decreases performance on SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%).
VQAv2 GQA Vizwiz SQA-I TextVQA POPE MME MMBench (en) MMBench (cn) LLAVA-W Rel. (%)
@@ D.2 Visual dependency influence ranking
Table 8: Cross-Architecture Generalization. Our LLaVA-ICONS-133K selected via LLaVA-v1.5-7B model (7B-selected) shows strong cross-architecture transferability, achieving 97.3% Rel., while the data selected via LLaVA-v1.5-13B model (13B-selected) reaches 98.1%, showing that our selected subset generalizes well to different architectures.
D.3 Cross-architecture generalization
@@
We further conduct experiments on cross architecture generalization to evaluate the transferability of our selected data across different model scales. While our subset was initially selected using LLaVA-v1.5-7B as the base model, we investigate whether these same examples remain effective for training larger architectures like LLaVA-v1.5-13B. This tests whether our selection criteria identify universally valuable training examples rather than model-specific patterns. Our results in Tab. 8 show cross-architecture generalization, with 13B model trained on 7B-selected subset achieving 98.1% Rel.. Both 7B-selected and 13B-selected subsets outperform random selection (95.7%), with the 13B-selected option showing particular strength in reasoning tasks like MMBench and POPE. This suggests that our selected subset captures fundamental visual-language understanding patterns that generalize well across different model architectures.